# Sliced-Wasserstein Estimation with Spherical Harmonics as Control Variates

### v1

@@ 1 Introduction
Informally, for $p\geq 1$ the $\operatorname{SW}_{p}$ distance between probability distributions $\mu$ and $\nu$ on $\mathbb{R}^{d}$ is defined as the integral value, over $\theta$ uniformly distributed on the unit sphere $\mathbb{S}^{d-1}$, of the Wasserstein distance $\mathrm{W}_{p}$ between the projections of distributions $\mu$ and $\nu$ on the direction of $\theta$. It takes advantage of the fact that (a) the Wasserstein distance between one-dimensional measures has a simple analytical form, being the $L^{p}([0,1])$-distance between their quantile functions (approximated using empirical quantiles), and (b) the integral is approximated by Monte Carlo sampling (uniformly on the unit sphere). As a consequence, two types of error affect the accuracy of most algorithms estimating the SW distance: first, the statistical error, due to the approximation of the distributions $\mu$ and $\nu$ by their empirical counterparts, which scales as $m^{-1/2}$, with $m$ the size of the samples observed from $\mu$ and $\nu$ (Nietert et al., 2022); second, the Monte Carlo integration error, which scales as $n^{-1/2}$, with $n$ the number of particles, as demonstrated in Nadjahi et al. (2020b) and Nietert et al. (2022). Importantly, these results establish that the sample complexity to approximate the SW distance does not depend on the dimension. As pointed out in Nadjahi et al. (2020b), even though the SW sample complexity is reduced compared to Wasserstein, it requires an additional Monte Carlo step which might offset the previous benefits if the associated variance is large.
@@ 2.2 Sliced-Wasserstein Distances
In practice, the SW distance of Equation 2 is approximated using a standard Monte Carlo estimate by (i) sampling a number $n\geq 1$ of independent random directions $\theta_{1},\ldots,\theta_{n}$ from $\mathrm{P}$ on $\mathbb{S}^{d-1}$ and (ii) by averaging the values $(f^{(p)}_{\mu,\nu}(\theta_{i}))_{i}$. This is described in Algorithm 1.
@@ 3 Monte Carlo with Control Variates
In practice, these quantities are estimated using the particles $\theta_{1},\ldots,\theta_{n}$ sampled from $\mathrm{P}$. The Ordinary Least Squares Monte Carlo (OLSMC) estimate $\mathrm{I}_{n}^{\mathrm{ols}}(f)$ is defined [^1] It is well-defined if and only if $\mathds{1}_{n}\notin\operatorname{Span}(\Phi)$. If ill-defined, one can reduce the number of control variates such that $s<n$. by
@@ 4.1 Spherical Harmonics
Note that, when restricted to the sphere, all polynomials are linear combinations of spherical harmonics. Moreover, spherical harmonics of different degrees are orthogonal with respect to $\langle f,g\rangle=\int_{\mathbb{S}^{d-1}}f(\theta)g(\theta)\,\mathrm{d}\mathrm{P}(\theta)$, where $\mathrm{P}$ is the uniform distribution on $\mathbb{S}^{d-1}$ (Dai, 2013, Theorem 1.1.2). In practice, orthonormal bases can be constructed by applying the Gram–Schmidt process and an explicit basis of spherical harmonics can be written in terms of the Gegenbauer polynomials in spherical coordinates (Dai, 2013, Theorem 1.5.1). In the following, let $\varphi_{\ell,k}$ with $\ell\geq 0$ and $k=1,\ldots,N_{\ell}^{d}$ denote a set of spherical harmonics forming an orthonormal basis of $\mathscr{H}_{\ell}^{d}$.
@@
The spherical harmonics $\{\varphi_{\ell,k},0\leq\ell,1\leq k\leq N_{\ell}^{d}\}$ form an orthonormal eigenbasis of the Hilbert space $L_{2}(\mathbb{S}^{d-1})$ so that for every $f\in L_{2}(\mathbb{S}^{d-1})$ we have
[…]
@@
$\mathrm{I}(\varphi_{\ell,k})=\int_{\mathbb{S}^{d-1}}\varphi_{\ell,k}(\theta)\,\mathrm{d}\mathrm{P}(\theta)=0.$ (7)
@@ 4.2 Spherical Harmonics Control Variates Estimate
For $\mu,\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$, the SHCV estimate of maximum degree $2L$ based on particles $\theta_{1},\ldots,\theta_{n}\in\mathbb{S}^{d-1}$ and control variates $\varphi=(\varphi_{j})_{j=1}^{s_{L,d}}$ is defined as the intercept of the OLS problem (6) with the integrand $f_{\mu,\nu}^{(p)}$ as response and all spherical harmonics of even degree from $2$ up to $2L$ as $n\times s_{L,d}$ covariate matrix $\Phi=(\varphi(\theta_{i})^{\top})_{i=1}^{n}$, i.e.,
@@
The procedure is described in Algorithm 2.
@@
4:  Return ${\color[rgb]{0.80078125,0.46875,0.73828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.46875,0.73828125}\pgfsys@color@rgb@stroke{0.80078125}{0.46875}{0.73828125}\pgfsys@color@rgb@fill{0.80078125}{0.46875}{0.73828125}\mathrm{SHCV}}_{n}=\mathrm{I}_{n}^{\mathrm{ols}}$
@@ 5 Numerical Experiments
To asses the finite-sample performance of the SHCV estimate, we first present some synthetic data examples involving the computation of $\operatorname{SW}_{2}^{2}(\mu,\nu)$ between multivariate Gaussian distributions. We consider different dimensions $d\in\{3;5;6;10;20\}$ and a varying number of random projections $n\in[10^{2};10^{4}]$. Then we focus on $\operatorname{SW}_{2}^{2}(\mu_{m},\nu_{m})$ between 3D point clouds from the ShapeNetCore dataset of Chang et al. (2015). In both cases, the ground truth of the Sliced-Wasserstein distance is estimated with a very large number $n=10^{8}$ of particles. The different results report the mean squared error (MSE) $\mathbb{E}[|{\operatorname{SW}}_{2,n}^{2}(\mu,\nu)-\operatorname{SW}_{2}^{2}(\mu,\nu)|^{2}]$ for the different estimates ${\operatorname{SW}}_{2,n}^{2}(\mu,\nu)$ where the expectation is computed as an average over $100$ independent runs. Finally, in the spirit of Kolouri et al. (2016) and Meunier et al. (2022), we implement a Kernel Support Vector Machine (SVM) classifier for image classification.
For reproducibility, the code is available upon request and numerical details are available in the appendices. In particular, we rely on the efficient implementation of Dutordoir et al. (2020) to build the spherical harmonics.
@@
Figure 2: MSE curves for Gaussian distributions in dimension $d\in\{3;6\}$ based on $n\in[10^{2};10^{4}]$ projections.
@@ 5.1 Synthetic Data: Multivariate Gaussians
We compute $\operatorname{SW}_{2}^{2}(\mu,\nu)$ between $\mu=\mathcal{N}(a,\mathbf{A})$ and $\nu=\mathcal{N}(b,\mathbf{B})$ with means $a,b\sim\mathcal{N}(\mathds{1}_{d},I_{d})$ and covariance matrices $\mathbf{A}=\Sigma_{a}\Sigma_{a}^{\top}$ and $\mathbf{B}=\Sigma_{b}\Sigma_{b}^{\top}$ where all the entries of $\Sigma_{a},\Sigma_{b}$ are drawn according to $\mathcal{N}(0,1)$. We consider here the exact case meaning that we directly evaluate the smooth integrand $f_{\mu,\nu}^{(2)}$ given in Remark 1. As the original lower-CV ($\text{CV}_{low}$) and upper-CV ($\text{CV}_{up}$) estimates are written to take discrete measures as inputs, we adapt them with the true means and covariance matrices. In Section B.2, we also consider the sampled case involving discrete measures $\mu_{m},\nu_{m}$ where we only have access to $m=1000$ samples $x_{1},\ldots,x_{m}\sim\mu$ and $y_{1},\ldots,y_{m}\sim\nu$.
Figure 2 reports the MSE of the different methods in dimension $d\in\{3;6\}$ with respect to the number of projections, while Figure 4 in Section B.1 does so with respect to computing time. Table 1 reports the MSE and computing times (in ms) in higher dimension $d\in\{5;10;20\}$. Figure 5 and Tables 4 and 5 in Section B.2 deal with the sampled case accordingly. In each case, the SHCV estimate gives the best MSE performance with respect to both the number of projections and computing time, beating other baselines by a factor up to $100$.
@@
MC $2.6$e-$2$ $06.5\pm 2.3$ $8.4$e-$3$ $06.7\pm 0.7$ $2.0$e-$2$ $06.5\pm 0.3$
$\text{CV}_{low}$ $4.7$e-$3$ $18.1\pm 4.5$ $3.0$e-$3$ $18.4\pm 4.3$ $4.4$e-$3$ $18.9\pm 3.3$
$\text{CV}_{up}$ $1.5$e-$2$ $20.1\pm 5.2$ $7.0$e-$3$ $16.8\pm 3.2$ $1.6$e-$2$ $14.6\pm 1.6$
CVNN $7.7$e-$3$ $45.4\pm 12\phantom{.}$ $1.0$e-$2$ $62.0\pm 14\phantom{.}$ $2.6$e-$2$ $73.6\pm 12\phantom{.}$
QMC $4.9$e-$3$ $22.0\pm 7.1$ $4.1$e-$3$ $42.3\pm 9.7$ $1.3$e-$2$ $62.2\pm 6.1$
RQMC $7.5$e-$4$ $23.6\pm 6.7$ $2.4$e-$3$ $52.4\pm 12\phantom{.}$ $1.4$e-$2$ $84.3\pm 14\phantom{.}$
SHCV $\mathbf{4.8}$e-$\mathbf{5}$ $15.6\pm 4.5$ $\mathbf{1.7}$e-$\mathbf{3}$ $10.4\pm 4.9$ $\mathbf{2.7}$e-$\mathbf{3}$ $11.2\pm 5.2$
@@ 6 Theoretical Analysis
This section gathers several theoretical properties of the novel SHCV estimate. We first state some elementary properties of the underlying integration rule then provide a convergence rate in probability of the integration error.
@@ 6.2 Asymptotic Error Bound
We provide a convergence rate in probability [^3] Notation $\mathcal{O}_{\mathbb{P}}(\cdot)$ means stochastic boundedness. of the integration error $\operatorname{SHCV}_{n,L}^{p}(\mu,\nu)-\operatorname{SW}_{p}^{p}(\mu,\nu)$ for general $\mu,\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ and $p\in[1,\infty)$, where $\operatorname{SW}_{p}(\mu,\nu)$ is the $\operatorname{SW}$ distance with respect to the uniform distribution on the sphere and $\operatorname{SHCV}_{n,L}^{p}(\mu,\nu)$ is our estimate in Definition 4 based on $n$ Monte Carlo particles on the sphere and using spherical harmonics of even degree up to $2L$ as control variates. The guarantee relies on two properties: first, a general convergence rate of the integration error of the OLSCV estimate in Portier & Segers (2019, Theorem 1), available in Section C.2, and second, a bound on the uniform approximation error of a Lipschitz function on the sphere by its orthogonal projection on the space of polynomials (Newman & Shapiro, 1964, Theorem 2). The latter result applies since the integrand $f_{\mu,\nu}^{(p)}$ in Equation 3 is always at least Lipschitz: by Han (2023, Theorem 2.4), we have
@@
The modulus of continuity of $f\in\mathscr{C}(\mathbb{S}^{d-1})$ is $\omega(h)=\sup\{|f(\theta)-f(\gamma)|:\rho(\theta,\gamma)\leq h\}$ for $h>0$, with $\rho(\theta,\gamma)=\arccos\langle\theta,\gamma\rangle$ the length of the arc connecting $\theta,\gamma\in\mathbb{S}^{d-1}$.
@@
There exists an absolute constant $A>0$ such that for each dimension $d\geq 2$, each function $f\in\mathscr{C}(\mathbb{S}^{d-1})$ and each degree $L$ there exists a polynomial $P_{f,L}$ on $\mathbb{S}^{d-1}$ of degree at most $L$ such that: $\sup_{\theta\in\mathbb{S}^{d-1}}|f(\theta)-P_{f,L}(\theta)|\leq A\,\omega(d/L).$
The polynomial $P_{f,L}$ in Theorem 1 that yields the uniform approximation is in general not the same as the orthogonal projection $\Pi_{f,L}=\sum_{\ell=0}^{L}\sum_{k=1}^{N_{\ell}^{d}}\hat{f}_{\ell,k}\varphi_{\ell,k}$ of $f$ in $L_{2}(\mathbb{S}^{d-1})$ on the space of spherical harmonics of degree at most $L$, where we use the notation of Proposition 1. Still, the $L_{2}(\mathbb{S}^{d-1})$ approximation error of $\Pi_{f,L}$ is bounded by the one of $P_{f,L}$, which is in turn bounded by the uniform approximation error of $P_{f,L}$. Applying this argument to the integrand $f_{\mu,\nu}^{(p)}$, we obtain the following convergence rate.
@@
Let $d\geq 2$, $p\in[1,\infty)$ and $\mu,\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ be fixed. For any degree sequence $L=L_{n}$ such that $L=o(n^{1/(2(d-1))})$ as $n\to\infty$, the integration error satisfies

### v2

@@ 1 Introduction
Informally, for $p\geq 1$ the $\operatorname{SW}_{p}$ distance between probability distributions $\mu$ and $\nu$ on $\mathbb{R}^{d}$ is defined as the integral value, over $\theta$ uniformly distributed on the unit sphere $\mathbb{S}^{d-1}$, of the Wasserstein distance $\mathrm{W}_{p}$ between the projections of distributions $\mu$ and $\nu$ on the direction of $\theta$. It takes advantage of the fact that (a) the Wasserstein distance between one-dimensional measures has a simple analytical form, being the $L^{p}([0,1])$-distance between their quantile functions (approximated using empirical quantiles), and (b) the integral is approximated by Monte Carlo sampling (uniformly on the unit sphere). As a consequence, two types of error affect the accuracy of most algorithms estimating the SW distance: first, the statistical error, due to the approximation of the distributions $\mu$ and $\nu$ by their empirical counterparts, which scales as $m^{-1/2}$, with $m$ the size of the samples observed from $\mu$ and $\nu$ (Nietert et al., 2022); second, the Monte Carlo integration error, which scales as $n^{-1/2}$, with $n$ the number of projections, as demonstrated in Nadjahi et al. (2020b) and Nietert et al. (2022). Importantly, these results establish that the sample complexity to approximate the SW distance does not depend on the dimension. As pointed out in Nadjahi et al. (2020b), even though the SW sample complexity is reduced compared to Wasserstein, it requires an additional Monte Carlo step which might offset the previous benefits if the associated variance is large.
@@ 2.2 Sliced-Wasserstein Distances
In practice, the SW distance of Equation 2 is approximated using a standard Monte Carlo estimate by (i) sampling a number $n\geq 1$ of independent random directions $\theta_{1},\ldots,\theta_{n}$ from $\mathrm{P}$ on $\mathbb{S}^{d-1}$ and (ii) by averaging the values $(f^{(p)}_{\mu,\nu}(\theta_{i}))_{i}$. This is described in Algorithm 1. In the paper, we will often work with the uniform distribution on the sphere, denoted by $\mathrm{U}$ or $\mathcal{U}(\mathbb{S}^{d-1})$.
@@ 3 Monte Carlo with Control Variates
In practice, these quantities are estimated using the directions $\theta_{1},\ldots,\theta_{n}$ sampled from $\mathrm{P}$. The Ordinary Least Squares Monte Carlo (OLSMC) estimate $\mathrm{I}_{n}^{\mathrm{ols}}(f)$ is [^1] It is well-defined if and only if $\mathds{1}_{n}\notin\operatorname{Span}(\Phi)$. If ill-defined, one can reduce the number of control variates such that $s<n$.
@@ 4.1 Spherical Harmonics
Note that, when restricted to the sphere, all polynomials are linear combinations of spherical harmonics. Moreover, spherical harmonics of different degrees are orthogonal with respect to $\langle f,g\rangle=\int_{\mathbb{S}^{d-1}}f(\theta)g(\theta)\,\mathrm{d}\mathrm{U}(\theta)$, where $\mathrm{U}$ is the uniform distribution on $\mathbb{S}^{d-1}$ (Dai, 2013, Theorem 1.1.2). In practice, orthonormal bases can be constructed by applying the Gram–Schmidt process and an explicit basis of spherical harmonics can be written in terms of the Gegenbauer polynomials in spherical coordinates (Dai, 2013, Theorem 1.5.1). In the following, let $\varphi_{\ell,k}$ with $\ell\geq 0$ and $k=1,\ldots,N_{\ell}^{d}$ denote a set of spherical harmonics forming an orthonormal basis of $\mathscr{H}_{\ell}^{d}$.
@@
The spherical harmonics $\{\varphi_{\ell,k},0\leq\ell,1\leq k\leq N_{\ell}^{d}\}$ form an orthonormal eigenbasis of the Hilbert space $L_{2}(\mathbb{S}^{d-1},\mathrm{U})$, so that for every $f\in L_{2}(\mathbb{S}^{d-1})$ we have
$f=\sum_{\ell=0}^{\infty}\sum_{k=1}^{N_{\ell}^{d}}\hat{f}_{\ell,k}\varphi_{\ell,k}\quad\text{where}\quad\hat{f}_{\ell,k}=\int f\varphi_{\ell,k}\,\mathrm{d}\mathrm{U}.$
@@
$\mathrm{I}(\varphi_{\ell,k})=\int_{\mathbb{S}^{d-1}}\varphi_{\ell,k}(\theta)\,\mathrm{d}\mathrm{U}(\theta)=0.$ (7)
@@ 4.2 Spherical Harmonics Control Variates Estimate
For $\mu,\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$, the SHCV estimate of maximum degree $2L$ based on projections $\theta_{1},\ldots,\theta_{n}\in\mathbb{S}^{d-1}$ and control variates $\varphi=(\varphi_{j})_{j=1}^{s_{L,d}}$ is defined as the intercept of the OLS problem (6) with the integrand $f_{\mu,\nu}^{(p)}$ as response and all spherical harmonics of even degree from $2$ up to $2L$ as $n\times s_{L,d}$ covariate matrix $\Phi=(\varphi(\theta_{i})^{\top})_{i=1}^{n}$, i.e.,
@@
The procedure is described in Algorithm 2. In the algorithm, the number of spherical harmonics is specified in advance. A selection based on a lasso procedure with cross-validation for the penalization parameter would be possible too, but is not pursued here for the sake of computing efficiency.
@@
4:  Return ${\color[rgb]{0.80078125,0.46875,0.73828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.80078125,0.46875,0.73828125}\mathrm{SHCV}}_{n}=\mathrm{I}_{n}^{\mathrm{ols}}$
@@ 5 Numerical Experiments
To asses the finite-sample performance of the SHCV estimate, we first present some synthetic data examples involving the computation of $\operatorname{SW}_{2}^{2}(\mu,\nu)$ between multivariate Gaussian distributions. We consider different dimensions $d\in\{3;5;6;10;20\}$ and a varying number of random projections $n\in[10^{2};10^{4}]$. Then we focus on $\operatorname{SW}_{2}^{2}(\mu_{m},\nu_{m})$ between 3D point clouds from the ShapeNetCore dataset of Chang et al. (2015). In both cases, the ground truth of the Sliced-Wasserstein distance is estimated with a very large number $n=10^{8}$ of projections. The different results report the mean squared error (MSE) $\mathbb{E}[|{{\operatorname{SW}}_{2,n}^{2}}(\mu,\nu)-\operatorname{SW}_{2}^{2}(\mu,\nu)|^{2}]$ for the different estimates ${{\operatorname{SW}}_{2,n}^{2}}(\mu,\nu)$ where the expectation is computed as an average over $100$ independent runs. Finally, in the spirit of Kolouri et al. (2016) and Meunier et al. (2022), we implement a Kernel Support Vector Machine (SVM) classifier for image classification. For ease of reproducibility, numerical details are in the appendix and the code is available here. In particular, we rely on the efficient implementation of Dutordoir et al. (2020) to build the spherical harmonics. The degrees of the spherical harmonics are reported in Table 9 in Section D.2.
@@
Figure 2: MSE curves for sampled Gaussian distributions in dimension $d\in\{3;6\}$.
@@ 5.1 Synthetic Data: Multivariate Gaussians
We compute $\operatorname{SW}_{2}^{2}(\mu_{m},\nu_{m})$ between empirical distributions $\mu_{m}=m^{-1}\sum_{i=1}^{m}\delta_{x_{i}}$ and $\nu_{m}=m^{-1}\sum_{j=1}^{m}\delta_{y_{j}}$ where $x_{1},\ldots,x_{m}\sim\mu=\mathcal{N}(a,\mathbf{A})$, $y_{1},\ldots,y_{m}\sim\nu=\mathcal{N}(b,\mathbf{B})$ with $m=1000$ samples. The parameters of the Gaussians are $a,b\sim\mathcal{N}(\mathds{1}_{d},I_{d})$ and covariance matrices $\mathbf{A}=\Sigma_{a}\Sigma_{a}^{\top}$ and $\mathbf{B}=\Sigma_{b}\Sigma_{b}^{\top}$ where all entries of $\Sigma_{a},\Sigma_{b}$ are drawn according to $\mathcal{N}(0,1)$. In Section B.1, we also consider the exact case where we directly evaluate the smooth integrand $f_{\mu,\nu}^{(2)}$ given in Remark 1. In that case, as the original lower-CV ($\text{CV}_{low}$) and upper-CV ($\text{CV}_{up}$) estimates are written to take discrete measures as inputs, we adapt them with the true means and covariance matrices.
Figure 2 reports the MSE of the different methods in dimension $d\in\{3;6\}$ with respect to the number of projections, while Figure 5 in Section B.2 does so with respect to computing time. Table 1 reports the MSE and computing times (in ms) in higher dimension $d\in\{5;10;20\}$. Figure 4 and Table 3 in Section B.1 deal with the exact case accordingly. In each case, the SHCV estimate gives the best MSE performance with respect to both the number of projections and computing time, beating other baselines by a factor up to $100$.
@@
MC $1.45$e-$3$ $81.1\pm 3.5$ $9.45$e-$4$ $80.7\pm 4.4$ $1.47$e-$3$ $81.1\pm 1.8$
$\text{CV}_{\text{low}}$ $2.67$e-$4$ $79.7\pm 1.1$ $3.45$e-$4$ $80.1\pm 1.4$ $3.82$e-$4$ $80.0\pm 1.0$
$\text{CV}_{\text{up}}$ $8.44$e-$4$ $83.0\pm 1.2$ $7.51$e-$4$ $83.0\pm 1.7$ $1.09$e-$3$ $83.1\pm 1.5$
CVNN $4.29$e-$4$ $110\phantom{.}\pm 2.2$ $1.12$e-$3$ $122\phantom{.}\pm 1.6$ $2.14$e-$3$ $127\phantom{.}\pm 1.4$
QMC $2.91$e-$4$ $100\phantom{.}\pm 1.2$ $2.37$e-$4$ $113\phantom{.}\pm 1.4$ $6.60$e-$4$ $129\phantom{.}\pm 1.4$
RQMC $5.80$e-$5$ $96.3\pm 2.2$ $2.75$e-$4$ $113\phantom{.}\pm 1.2$ $1.17$e-$3$ $130\phantom{.}\pm 1.0$
SHCV $\mathbf{2.68}$e-$\mathbf{6}$ $89.0\pm 6.3$ $\mathbf{1.93}$e-$\mathbf{4}$ $89.0\pm 4.5$ $\mathbf{2.95}$e-$\mathbf{4}$ $88.1\pm 2.8$
@@ 6 Theoretical Analysis
This section gathers several theoretical properties of the novel SHCV estimate. We first state some elementary properties of the underlying integration rule with respect to a general distribution $\mathrm{P}$ on the sphere and then provide a convergence rate in probability of the integration error in case of the uniform distribution $\mathrm{U}$ on the sphere.
@@ 6.2 Asymptotic Error Bound
We provide a convergence rate in probability [^3] The notation $R_{n}=\mathcal{O}_{\mathbb{P}}(a_{n})$, for random variables $R_{n}$ and constants $a_{n}>0$, means that for every $\delta>0$ there exists $K_{\delta}>0$ such that $\mathbb{P}(|R_{n}|>K_{\delta}a_{n})<\delta$ for all $n$.of the integration error $\operatorname{SHCV}_{n,L}^{p}(\mu,\nu)-\operatorname{SW}_{p}^{p}(\mu,\nu)$ for general probability measures $\mu,\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ and $p\in[1,\infty)$, where $\operatorname{SW}_{p}(\mu,\nu)$ is the $\operatorname{SW}$ distance with respect to the uniform distribution on the sphere and $\operatorname{SHCV}_{n,L}^{p}(\mu,\nu)$ is our estimate in Definition 4 based on $n$ random projections on the sphere and using spherical harmonics of even degree up to $2L$ as control variates. The guarantee relies on two properties: first, a general convergence rate of the integration error of the OLSCV estimate in Portier & Segers (2019, Theorem 1), available in Section C.2, and second, a bound on the uniform approximation error of a Lipschitz function on the sphere by its orthogonal projection on the space of polynomials (Newman & Shapiro, 1964, Theorem 2). The latter result applies since the integrand $f_{\mu,\nu}^{(p)}$ in Equation 3 is always at least Lipschitz: by Han (2023, Theorem 2.4), we have
@@
The modulus of continuity of $f\in\mathscr{C}(\mathbb{S}^{d-1})$ is defined for $h>0$ by
$\omega(h)=\sup\{|f(\theta)-f(\gamma)|:\rho(\theta,\gamma)\leq h\}$
with $\rho(\theta,\gamma)=\arccos\langle\theta,\gamma\rangle$ the length of the arc connecting $\theta,\gamma\in\mathbb{S}^{d-1}$.
@@
There exists an absolute constant $A>0$ such that for each dimension $d\geq 2$, each function $f\in\mathscr{C}(\mathbb{S}^{d-1})$ and each degree $L$ there exists a polynomial $P_{f,L}$ on $\mathbb{S}^{d-1}$ of degree at most $L$ such that
$\sup_{\theta\in\mathbb{S}^{d-1}}|f(\theta)-P_{f,L}(\theta)|\leq A\,\omega(d/L).$
The polynomial $P_{f,L}$ in Theorem 1 that yields the uniform approximation is in general not the same as the orthogonal projection $\Pi_{f,L}=\sum_{\ell=0}^{L}\sum_{k=1}^{N_{\ell}^{d}}\hat{f}_{\ell,k}\varphi_{\ell,k}$ of $f$ in $L_{2}(\mathbb{S}^{d-1})$ on the space of spherical harmonics of degree at most $L$, where we use the notation of Proposition 1. Still, the $L_{2}(\mathbb{S}^{d-1})$ approximation error of $\Pi_{f,L}$ is bounded by the one of $P_{f,L}$, which is in turn bounded by the uniform approximation error of $P_{f,L}$. Applying this argument to the integrand $f_{\mu,\nu}^{(p)}$, we obtain the following convergence rate in probability of the SHCV estimate.
@@
Let $d\geq 2$, $p\in[1,\infty)$ and $\mu,\nu\in\mathcal{P}_{p}(\mathbb{R}^{d})$ be fixed. For any degree sequence $L=L_{n}$ such that $L=o(n^{1/(2(d-1))})$ as $n\to\infty$, the integration error of the SHCV estimate with respect to the uniform distribution on the sphere satisfies